# Challenges in numerical integration in physics-informed neural networks modelling

### v2

@@ 3.3 Calculation of Fourier coefficients
integral 1000 phi_double -0.132240670973319 0.178 12 Yes1
phi_vpa -0.132240669593892 0.198 1 No
phi_exact -0.132240669593892 0.080 1 No
quad 1000 phi_double -0.132240607546634 0.225 30 No
phi_vpa -0.132240616544284 0.438 30 No
phi_exact -0.132240616544284 0.225 30 No
quadl 1000 phi_double -0.132240665269117 0.107 8 No
phi_vpa -0.132240669544252 0.185 8 No
phi_exact -0.132240669544252 0.105 8 No
trapz 1000 phi_double -0.132240669210701 0.063 1 No
phi_vpa -0.132240669593892 5.482 1 No
phi_exact -0.132240669593892 0.061 1 No
integral 10000 phi_double -0.132241272405812 0.190 11 Yes1
phi_vpa -0.132240669593892 0.181 1 No
phi_exact -0.132240669593892 0.078 1 No
quad 10000 phi_double -0.132263860323737 0.272 38 No
phi_vpa -0.132240616544284 0.441 30 No
phi_exact -0.132240616544284 0.223 30 No
quadl 10000 phi_double -0.132218822360428 0.109 8 No
phi_vpa -0.132240669544252 0.189 8 No
phi_exact -0.132240669544252 0.105 8 No
trapz 10000 phi_double -0.132242712125416 0.060 1 No
phi_vpa -0.132240669593892 5.652 1 No
phi_exact -0.132240669593892 0.069 1 No
integral 100000 phi_double -0.164105328139715 0.212 12 Yes1
phi_vpa -0.132240669593892 0.193 1 No
phi_exact -0.132240669593892 0.083 1 No
quad 100000 phi_double -0.217677705429780 27.429 5006 Yes2
phi_vpa -0.132240616544284 0.505 30 No
phi_exact -0.132240616544284 0.257 30 No
quadl 100000 phi_double -0.175317505741447 11.336 2012 Yes2
phi_vpa -0.132240669544252 0.189 8 No
phi_exact -0.132240669544252 0.110 8 No
trapz 100000 phi_double -0.162243580664938 0.060 1 No
phi_vpa -0.132240669593892 5.521 1 No
phi_exact -0.132240669593892 0.063 1 No
integral 250000 phi_double 0.000000000000000 0.080 1 No
phi_vpa -0.132240669593892 0.245 1 No
phi_exact -0.132240669593892 0.076 1 No
quad 250000 phi_double 0.000000000000000 0.081 4 No
phi_vpa -0.132240616544284 0.436 30 No
phi_exact -0.132240616544284 0.225 30 No
quadl 250000 phi_double 0.000000000000000 0.073 2 No
phi_vpa -0.132240669544252 0.191 8 No
phi_exact -0.132240669544252 0.108 8 No
trapz 250000 phi_double 0.000000000000000 0.060 1 No
phi_vpa -0.132240669593892 5.608 1 No
phi_exact -0.132240669593892 0.062 1 No
4 Heat equation problem
@@
$\dfrac{\partial u}{\partial t}(t,x)-\alpha\dfrac{\partial^{2}u}{\partial x^{2}}(t,x)$ $=0,\quad t\in(0,1],x\in(-1,1),$ (10)
$u(t,-1)=u(t,1)$ $=0,\quad t\in[0,1],$ (11)
$u(0,x)$ $={\color[rgb]{0,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.5}\phi(x)},\quad x\in[-1,1].$ (12)
Solution by separation of variables with $f_{n}(t)=\mathrm{e}^{-\alpha t((2n-1)\pi/2)^{2}}$, base functions $g_{n}(x)$ given by (8) and Fourier coefficients $A_{n}$ given by (9) takes the form
@@
5 Physics-informed neural network model
@@
We now train a physics informed neural network (PINN) to numerically compute the solution of the heat equation by using the limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) algorithm.
@@
    • Multilayer perceptron DNN architecture with 9 fully connect hidden layers with 20 neurons each.
@@
    • As activation functions we take hyperbolic tangents (for demonstration purposes).
@@ 6 Conclusion
Calculating the value of a function is a basic numerical operation in most numerical algorithms. Often, we incorrectly automatically assume that the value calculated is within machine precision in a given arithmetic, not realizing that this may not be the case. Inaccurate function values can subsequently cause very unreliable inputs to algorithms working with that function. Yet we may not even notice that the obtained results are incorrect.
In this article, we have shown examples of how inaccurate function values can negatively influence the results of numerical integration (quadrature) and subsequently modelling using PINNs.
For calculations, we are interested in the accuracy of the result and the computational time. Requested optimal combination is to have an accurate result in minimum time. We have shown in the examples that a longer computational time does not automatically mean a more accurate result, on the contrary, it can cause problems. On the other hand, a sufficiently accurate result may not always be time consuming. It is worth to mention that one has to be careful with low computational times that can sometimes occur for inaccurately evaluated integrand. Therefore, for each computation, it is important to analyze the entire algorithm including all the data.
In contrast to the results of Daněk and Pospíšil (2020), where authors propose a regime switching algorithm to change the integrand evaluation to vpa only in problematic cases, here the PDE solution method is significantly improved only by evaluating the initial condition in vpa.
@@
Numerical integration challenges, as discussed in this paper, arise due to inaccuracies in evaluating integrands, leading to severe precision problems in many models, and as demonstrated in particular in physics-informed neural networks. These issues stem from floating-point arithmetic limitations and improper numerical quadrature methods, especially when model parameters exacerbate these problems. Similar challenges can occur for example in the peridynamic models, which rely on integral formulations to capture long-range interactions and avoid spatial derivatives. The evaluation of integro-differential operators in peridynamics, as seen in its bond-based models, requires precise numerical integration for accurate representation of material behavior, including crack propagation and heterogeneous material response (Haghighat, Bekar, Madenci, and Juanes, 2021; Ning, Cai, Dong, Liu, and Wang, 2023).

### v3

@@ 3.3 Calculation of Fourier coefficients
integral 1000 phi_double -0.132240670973319 0.117 12 Yes1
phi_vpa -0.132240669593892 0.075 1 No
phi_exact -0.132240669593892 0.043 1 No
quad 1000 phi_double -0.132240607546634 0.106 30 No
phi_vpa -0.132240616544284 0.203 30 No
phi_exact -0.132240616544284 0.108 30 No
quadl 1000 phi_double -0.132240665269117 0.058 8 No
phi_vpa -0.132240669544252 0.094 8 No
phi_exact -0.132240669544252 0.060 8 No
trapz 1000 phi_double -0.132240669210701 0.034 1 No
phi_vpa -0.132240669593892 2.396 1 No
phi_exact -0.132240669593892 0.035 1 No
tanh-sinh 1000 phi_double -0.132240685879949 0.033 1 No
phi_vpa -0.132240671735414 0.080 1 No
phi_exact -0.132240671735414 0.033 1 No
integral 10000 phi_double -0.132241272405812 0.091 11 Yes1
phi_vpa -0.132240669593892 0.074 1 No
phi_exact -0.132240669593892 0.046 1 No
quad 10000 phi_double -0.132263860323737 0.129 38 No
phi_vpa -0.132240616544284 0.199 30 No
phi_exact -0.132240616544284 0.105 30 No
quadl 10000 phi_double -0.132218822360428 0.053 8 No
phi_vpa -0.132240669544252 0.092 8 No
phi_exact -0.132240669544252 0.054 8 No
trapz 10000 phi_double -0.132242712125416 0.039 1 No
phi_vpa -0.132240669593892 2.380 1 No
phi_exact -0.132240669593892 0.037 1 No
tanh-sinh 10000 phi_double -0.132181727630471 0.033 1 No
phi_vpa -0.132240671735414 0.050 1 No
phi_exact -0.132240671735414 0.032 1 No
integral 100000 phi_double -0.164105328139715 0.124 12 Yes1
phi_vpa -0.132240669593892 0.070 1 No
phi_exact -0.132240669593892 0.043 1 No
quad 100000 phi_double -0.217677705429780 11.343 5006 Yes2
phi_vpa -0.132240616544284 0.196 30 No
phi_exact -0.132240616544284 0.108 30 No
quadl 100000 phi_double -0.175317505741447 4.736 2012 Yes2
phi_vpa -0.132240669544252 0.094 8 No
phi_exact -0.132240669544252 0.056 8 No
trapz 100000 phi_double -0.162243580664938 0.037 1 No
phi_vpa -0.132240669593892 2.380 1 No
phi_exact -0.132240669593892 0.033 1 No
tanh-sinh 100000 phi_double -0.114446017197740 0.032 1 No
phi_vpa -0.132240671735414 0.053 1 No
phi_exact -0.132240671735414 0.060 1 No
integral 250000 phi_double 0.000000000000000 0.043 1 No
phi_vpa -0.132240669593892 0.071 1 No
phi_exact -0.132240669593892 0.045 1 No
quad 250000 phi_double 0.000000000000000 0.043 4 No
phi_vpa -0.132240616544284 0.204 30 No
phi_exact -0.132240616544284 0.105 30 No
quadl 250000 phi_double 0.000000000000000 0.039 2 No
phi_vpa -0.132240669544252 0.094 8 No
phi_exact -0.132240669544252 0.054 8 No
trapz 250000 phi_double 0.000000000000000 0.035 1 No
phi_vpa -0.132240669593892 2.393 1 No
phi_exact -0.132240669593892 0.033 1 No
tanh-sinh 250000 phi_double 0.000000000000000 0.033 1 No
phi_vpa -0.132240671735414 0.054 1 No
phi_exact -0.132240671735414 0.033 1 No
4 Problematic use-cases
In this section we introduce three problematic use-cases, one linear and one non-linear ordinary differential equation (ODE) and one linear partial differential equation (PDE).
4.1 Linear ODE
We consider an initial problem for linear ODE of the second order
$x^{\prime\prime}(t)+k^{2}x(t)=0,\quad x(0)=0,\,x^{\prime}(0)=k,$ (10)
where $k\in\mathbf{R}$ is a parameter. Initial value problem (10) has the unique exact solution $x(t)=\sin(k\cdot t)$.
Let $R(t)=x^{\prime\prime}(t)+k^{2}x(t)$ be the residuum for problem (10) and let us calculate it using the automatic differentiation over the first three periods of the solution. In Figure 7 we can see that the increasing value of $k$ causes also the higher value of the residua that should be theoretically zero. This problem belongs to the category of highly oscillatory functions mentioned in Section 2.1.
Figure 7: Residuum $R(t)$ calculated using AD for different values of $k$.
4.2 Non-linear ODE
We consider an initial problem for the van der Pol oscillator modelled as the non-linear ODE of the second order
$x^{\prime\prime}(t)-\mu(1-x^{2}(t))x^{\prime}(t)+x(t)=0,\quad x(0)=2,\,x^{\prime}(0)=0,$ (11)
where $\mu\in\mathbf{R}$ is a dumping parameter. Although the theorem of Picard–Lindelöf guarantees a unique analytic solution, it does not have an exact solution in closed form, it can only be written as a Taylor (Maclaurin) series. We can rewrite (11) as the system
$\begin{split}x^{\prime}(t)&=y(t),\\ y^{\prime}(t)&=\mu(1-x^{2}(t))y(t)-x(t),\\ x(0)&=2,\\ y(0)&=0.\end{split}$ (12)
To solve the system (12) numerically, we may use some suitable solver for stiff differential equations such as ode15s in MATLAB. In Section 5.2 we take ode15s solution as a reference when comparing it to a solution using PINNs.
4.3 Heat equation problem
@@
$\begin{split}\dfrac{\partial u}{\partial t}(t,x)-\alpha\dfrac{\partial^{2}u}{\partial x^{2}}(t,x)&=0,\quad t\in(0,1],x\in(-1,1),\\ u(t,-1)=u(t,1)&=0,\quad t\in[0,1],\\ u(0,x)&={\color[rgb]{0,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.5}\phi(x)},\quad x\in[-1,1].\end{split}$ (13)
A solution by separation of variables with $f_{n}(t)=\mathrm{e}^{-\alpha t((2n-1)\pi/2)^{2}}$, base functions $g_{n}(x)$ given by (8) and Fourier coefficients $A_{n}$ given by (9) takes the form
@@
5 Physics-informed neural network models
In this section we now solve all three problematic cases from Section 4 using PINNs.
5.1 Linear ODE
Let us now consider a PINNs solution of the problem (4.1). Training data are generated on the go in $N_{f}=128$ random collocation points $t_{i}$ covering the first period, i.e. the interval $[0,2\pi/k]$. We use a multilayer perceptron NN architecture with 2 fully connected hidden layers with 50 neurons each. Input layer has one input channel corresponding to the input $t$, output layer has one output channel $x(t)$. As activation functions, we take hyperbolic tangents. Total number of learnable parameters is 2700, all of type double. During the NN training, the Adam optimizer is used with the default setting (beta1=0.9, beta2=0.999 and epsilon=1e-8). In the following experiments, number of epochs and the value of the learning rate (LR) differs.
We consider the model loss function:
$\text{loss}=\text{MSE}_{R}+\lambda\cdot\text{MSE}_{I},\text{ where }\text{MSE}_{R}=\frac{1}{N_{f}}\sum\limits_{i=1}^{N_{f}}\left|R(t_{i})\right|^{2}\text{ and }\text{MSE}_{I}=x^{2}(0)+(x^{\prime}(0)-k)^{2}.$
Parameter $\lambda$ is used to put more stress on satisfying the initial conditions, in particular the value $\lambda=k^{2}$ is used. Calculating $\text{MSE}_{R}$ and $\text{MSE}_{I}$ requires the derivatives $x^{\prime}$ and $x^{\prime\prime}$ of the output $x$ of the model using automatic differentiation.
In Figure 10 we present a comparison of the PINNs solution to the exact solution for relatively small values of the parameter $k\in\{1,10,100,1000\}$. An increasing $k$ requires an increased number of epochs. The learning rate is changed adaptively (in a prescribed schedule) according to the Table 5 below. In all four experiments, the PINNs solution is satisfactory, revealing no numerical problems.
Table 5: Settings used in experiments depicted in Figure 10
$k$ 1 10 100 1000
number of epochs 1000 1000 3000 10000
LRinit 1e-2 1e-2 1e-2 1.3e-2
adaptive LR factor 0.99 0.99 0.98 0.97
after each # epoch 10 10 100 100
Figure 10: Comparison of the PINNs solution to the exact solution of problem (10), four experiments considered, see Table 5.
In Section 4.1 we showed that error of automatic differentiation applied to (10) increases with increasing value of $k$. In Figures 11 and 12 we consider only the value $k=1000$ that can lead to severe numerical problems of the PINNs solution when the setting is not carefully and suitably chosen. In particular, in Figure 11, only the initial learning rate is different compared to the settings $k=1000$ in Table 5 (the last column). Whereas on the left, the PINNs solution is biased with smaller error in initial conditions, the PINNs solution on the right is not satisfactory at all, in both cases after 10000 epochs. For convenience, we depict here also the learning rate decay as described in Table 5. Similar misbehaviour can be observed in the case, where the learning rate is fixed in all epochs, see Figure 12. Constant learning rate causes higher oscillations in the loss function (even with a big “jump” after ca 7500 epoch in the left case) and consequently worse results.
Figure 11: Comparison of the PINNs solution to the exact solution of problem (10), with $k=1000$ and adaptive LR with LRinit=1e-2 (left) and LRinit=1e-4 (right).
Figure 12: Comparison of the PINNs solution to the exact solution of problem (10), with $k=1000$ and fixed LR=1e-2 (left) and LR=1e-4 (right).
5.2 Nonlinear ODE
Let us now consider a PINNs solution of the problem (4.2). Training data are generated on the go in $N_{f}=128$ random collocation points $t_{i}$ covering the interval [0,7][0,7][ 0 , 7 ]. We use a multilayer perceptron NN architecture as in the previous section, i.e. with 2 fully connected hidden layers with 50 neurons each. Input layer has one input channel corresponding to the input $t$, output layer has one output channel $x(t)$. As activation functions, we take hyperbolic tangents. Total number of learnable parameters is again 2700, all of type double. During the NN training, the Adam optimizer is used with the default setting (beta1=0.9, beta2=0.999 and epsilon=1e-8). The learning rate is changed adaptively (in a prescribed schedule) according to the Table 6 below.
Table 6: Settings used in experiments depicted in Figures 13 and 14
$\mu$ 1 2 2 2 2 2
figure 13(a) 13(b) 14(a) 14(b) 14(c) 14(d)
number of epochs 3000 25000 50000 70000 50000 50000
LRinit 7e-3 8e-3 2.5e-2 6e-3 3e-3 5e-3
adaptive LR factor 0.97 0.99 0.99 0.99 0.99 0.99
after each # epoch 100 100 100 100 200 200
We consider the model loss function: $\text{loss}=\text{MSE}_{R}+\lambda\cdot\text{MSE}_{I}$, where
$\text{MSE}_{R}$ $=\frac{1}{N_{f}}\sum\limits_{i=1}^{N_{f}}\left|x^{\prime\prime}(t_{i})-\mu(1-x^{2}(t_{i}))x^{\prime}(t_{i})+x(t_{i})\right|^{2},$
$\text{MSE}_{I}$ $=(x(0)-2)^{2}+(x^{\prime}(0))^{2}.$
Parameter $\lambda$ can be used to put more stress on satisfying the initial conditions, however, in the following experiments when $\mu\in\{1,2\}$, the value of $\lambda$ does not bring any improvements, so that the value $\lambda=\mu$ is used for convenience. Calculating $\text{MSE}_{R}$ and $\text{MSE}_{I}$ requires the derivatives $x^{\prime}$ and $x^{\prime\prime}$ of the output $x$ of the model using automatic differentiation.
In Figure 13 we present a comparison of the PINNs solution to the reference numerical solution obtained by ode15s for different values of the parameter $\mu\in\{1,2\}$. For $\mu=1$ the PINNs solution is satisfactory almost regardless of the LR settings. For $\mu=2$, the PINNs solution becomes highly sensitive not only to LR settings, but also to a particular realization of the random number generator used for the stochastic gradient descent algorithm. In case (b), the PINNs solution is satisfactory, but it was obtained as one of the many realizations, although using the same LR settings.
  (a) (b)
Figure 13: Comparison of the PINNs solution (correct) to the reference numerical solution (ode15s) of problem (11) for (a) $\mu=1$ and (b) $\mu=2$.
In Figure 14 we give examples of non-correct PINNs solutions in comparison to the reference numerical solution obtained by ode15s for parameter $\mu=2$. In all four cases, the loss function is very small, but the solutions are far from being satisfactory. In cases (a) and (b), although we have a satisfactory overlap of the PINNs and the reference solution at the beginning of the time interval, the PINNs solution suddenly switches to the almost zero singular solution. Moreover, in case (c), the solution does not continue as a zero solution, and in case (d) the solution is completely wrong from the beginning of the considered time interval even after 50 thousand epochs.
  (a) (b)
Refer to caption Refer to caption
  (c) (d)
Figure 14: The examples of the PINNs solution (non-correct) to the reference numerical solution (ode15s) of problem (11) for $\mu=2$.
5.3 Heat equation problem
@@
We now train a PINN to numerically compute the solution of the heat equation (13) by using the limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) algorithm.
@@
    • Multilayer perceptron DNN architecture with 9 fully connected hidden layers with 20 neurons each.
@@
    • As activation functions we take hyperbolic tangents.
@@ 6 Conclusion
Calculating the value of a function together with its derivatives or integrals are basic numerical operations in most numerical algorithms. Often, we incorrectly automatically assume that the value calculated is within machine precision in a given arithmetic, not realizing that this may not be the case. Inaccurate function values can subsequently cause very unreliable inputs to algorithms working with that function. Yet we may not even notice that the obtained results are incorrect.
Automatic differentiation is a core tool to automate the simultaneous computation of the numerical values of arbitrarily complex functions and their derivatives. It computes exact, machine-precision derivatives by threading the chain rule through each operation as code runs. Unlike finite differences it needs no step-size tuning, and unlike symbolic math it acts on program instructions, avoiding complicated expression manipulations. Gradients and Jacobians emerge alongside function evaluation in forward or reverse mode, yet they still inherit floating-point round-off issues. It is worth to mention that neural network back-propagation is simply the reverse-mode automatic differentiation (Baydin, Pearlmutter, Radul, and Siskind, 2018) and hence many neural networks trainings can suffer from numerical precision if not treated correctly.
In this article, we have analysed three problematic use-cases of using PINNs to solve differential equations. In particular, a linear and a non-linear stiff ODE were considered to show how changing one model parameter or NN settings produces non-satisfactory results. To demonstrate the detection difficulty, we presented several experiments with very low loss function, but with the solution being highly biased. For a linear partial differential equation we show how inaccurate function values can negatively influence the results of numerical integration (quadrature) and subsequently the PINNs solution.
For quadrature calculations, we are interested in the accuracy of the result and the computational time. Requested optimal combination is to have an accurate result in minimum computational time. We have shown in the examples that a longer computational time does not automatically mean a more accurate result, on the contrary, it can cause problems. On the other hand, a sufficiently accurate result may not always be time consuming. It is worth mentioning that one has to be careful with low computational times that can sometimes occur for inaccurately evaluated integrand. Moreover, in practice, an integrand could be available only as a black box and hence it might be difficult to conclude that the problems are caused only by the inaccurately evaluated integrand. Therefore, for each computation, it is important to analyse the entire algorithm including all the data. In contrast to the results of Daněk and Pospíšil (2020), where authors propose a regime switching algorithm to change the integrand evaluation to vpa only in problematic cases, here the PDE solution method is significantly improved only by evaluating the initial condition in vpa.
@@
Numerical integration challenges, as discussed in this paper, arise due to inaccuracies in evaluating integrands, leading to severe precision problems in many models, and as demonstrated in particular in physics-informed neural networks. These issues stem from floating-point arithmetic limitations and improper numerical quadrature methods, especially when model parameters exacerbate these problems. Similar challenges can occur for example in the peridynamic models, which rely on integral formulations to capture long-range interactions and avoid spatial derivatives. The evaluation of integro-differential operators in peridynamics, as seen in its bond-based models, requires precise numerical integration for accurate representation of material behaviour, including crack propagation and heterogeneous material response (Haghighat, Bekar, Madenci, and Juanes, 2021; Ning, Cai, Dong, Liu, and Wang, 2023).